### 08 Jan 2024 v1

@@ Abstract
Fine-tuning studies whether some physical parameters, or relevant ratios between them, are located within so-called life-permitting intervals of small probability outside of which carbon-based life would not be possible. Recent developments have found estimates of these probabilities that circumvent previous concerns of measurability and selection bias. However, the question remains if fine-tuning can indeed be known. Using a mathematization of the epistemological concepts of learning and knowledge acquisition, we argue that most examples that have been touted as fine-tuned cannot be formally assessed as such. Nevertheless, fine-tuning can be known when the physical parameter is seen as a random variable and it is supported in the nonnegative real line, provided the size of the life-permitting interval is small in relation to the observed value of the parameter.
keywords:
active information, knowledge acquisition, learning
@@ 1 Introduction
Cosmological fine-tuning (FT) says that some physical parameters, some ratios between them, and some boundary conditions, must pertain to intervals of small probability to permit the existence of carbon-based life [1]. FT started to spark the interest of the scientific community when Carter [2] brought it to light calling it the anthropic principle, a term that, according to Lewis and Barnes [1], was popularized and modified by Barrow and Tipler [3]. Since then, considerable effort has been made to determine the length of these so-called life-permitting intervals (LPIs), constituting the area where the scientific literature has primarily focused [see, e.g., 4, 5, 6]. For this reason, some have linked FT with the small length of the intervals, with no regard for their probability [see, e.g., 7]. This evokes the idea of “naturalness” in physics, where a good theory is assumed to present numbers that are close to unity, and departures from it point to ad-hoc explanations trying to fit observations to theory as Mercury epicycles in the Ptolemaic model [8]. FT goes beyond naturalness, asserting that, according to our current theories, such unnaturalness is needed if carbon-based life is going to exist [9]. The existence of life sets a specification; i.e., a subset of possible outcomes of parameters for which life is permitted. These outcomes maximize a function $f$ that quantifies how specified outcomes are, and it is stochastically independent of the original random variable. In cosmological FT, this random variable is the physical parameter, whose value is assumed to be independent of the requisites for carbon-based life existence, and the LPI maximizes an indicator function $f$ that determines if life is possible or not. Therefore, the tuning of a physical constant for life is fine if and only if the probability of its LPI is small [10].
@@
Fine-tuning divides the scientific and philosophical communities between those who dismiss it as mere speculation [e.g., 8, 11, 12] and those who consider it a serious scientific endeavor [e.g., 13, 1, 14]. One of the main criticisms, even among those who see it as a legitimate question, is the lack of a valid probability distribution to impose over the parameters [4]. In this direction, Hossenfelder asks: “[H]ow do you make a statement about probability without a probability distribution?” [8, p. 205]. The problem is indeed daunting: it requires estimating the probability of the LPI using a biased sample of size 1 taken from an unknown distribution supported in an unknown space. More explicitly, and to establish some notation, let $\mathcal{X}$ be the possible values of a random physical parameter $X$. FT is about inferring the probability of its LPI $\ell_{X}$ when the only observation we possess is $x_{0}$, the observed value of $X$ in our universe; moreover, such an observation suffers from selection bias because that universe harbors life; and the only information about the probability distribution $F=F_{X}$ of $X$ is that $F$ is supported on $\mathcal{X}$ and $x_{0}\in\mathcal{X}$. Moreover, assuming Carter’s definition, it is just natural to assume that $F$ corresponds to a prior distribution of $X$ that maximizes the entropy [2].
@@
Past efforts to find the probabilities of LPIs use mainly uniform distributions over the set of possible values that constants might take [see, e.g., 15, 16, 17]. Nonetheless, this approach attests to the difficulty of the task, since the choice of the uniform distribution is based on Bernoulli’s Principle of Insufficient Reason [18], which requires knowledge that the sample space is finite —a strong assumption that cannot be warranted for cosmological FT. This legitimate criticism of FT measurements came from philosophy, where it was called the normalization problem [11, 12, 19, 20].
@@
Recently Díaz-Pachón, Hössjer, and Marks developed Algorithm 1 (below) to find an upper bound on the probability of the known LPI $\ell_{X}$ of a parameter $X$ [21, 22]. Algorithm 1 can be viewed as a statistical test, where the null hypothesis $H_{0}$ that tuning of $X$ is coarse is tested against the alternative hypothesis $H_{1}$ that $X$ is fine-tuned. The first two steps of Algorithm 1 ask for the input: a parameter space $\mathcal{X}$ (e.g., $\mathbb{N}$, $\mathbb{R}^{+}$, $\mathbb{R}$, $\mathbb{R}^{n}$, $\mathbb{C}$, or some subset thereof), and the constraints to impose on the prior distributions $F_{X}$ supported on $\mathcal{X}$ (e.g., if some event $B\subset\mathcal{X}$ has a known probability, if some of the moments of $X$ are finite, etc.). In the third step, a collection
@@
is obtained. In the fifth step, the outcome of Algorithm 1 is produced: if $\text{TP}_{\max}$ is small ($\leq\delta$ according to Definition 1) we reject the null hypothesis $H_{0}$ and conclude there is FT. If the maximum tuning probability is not small, the algorithm is inconclusive. In this case, we cannot reject $H_{0}$. This does not mean that $X$ is not finely tuned though. Indeed, $H_{1}$ may still be true when $H_{0}$ is not rejected, since $F_{0}(\ell_{X})$ may be larger of smaller than $\delta$ when $\text{TP}_{\max}>\delta$.
Algorithm 1 Algorithm for testing fine-tuning
1:Input: Choose the set of possible values $\mathcal{X}$ of $X$.
2:Input: Define constraints on the distribution $F=F_{X}$ of $X$.
@@
3:Find the family =${\cal F}=\{F(\cdot;\theta);\theta\in\Theta\}$ of maxent distributions $F$ with support $\mathcal{X}$, subject to the constraints in Step 2 and assume ($H_{0}$) that the true distribution of $X$ satisfies $F_{0}=F(\cdot;\theta_{0})\in\mathcal{F}$ for some $\theta_{0}\in\Theta$.
@@
4:Find the maximum tuning probability $\text{TP}_{\text{max}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\max\{F(\ell_{X});\,F\in{\cal F}\}$ over $\mathcal{F}$ for the life-permitting interval $\ell_{X}$.
@@
5:Output: If $\text{TP}_{\text{max}}$ is small ($\leq\delta$) there is fine-tuning (reject $H_{0}$); otherwise, the test is inconclusive (do not reject $H_{0}$).
@@
Algorithm 1 has some important properties. First, the upper bound $\text{TP}_{\text{max}}$ is sharp because if the family ${\cal F}$ is made of a single maxent distribution $F_{0}$, then $\text{TP}_{\max}=F_{0}(\ell_{X})$. For instance, assume that Step 1 selects $\mathcal{X}=\mathbb{R}^{+}$, and in Step 2 it is assumed that the first moment is finite and known (e.g., ${\mathbf{E}}_{0}X=\theta_{0}$). Then the family of Step 3 is made of a single distribution with density $F^{\prime}(x;\theta_{0})=f(x;\theta_{0})=e^{-x/\theta_{0}}/\theta_{0}$, the exponential with mean $\theta_{0}$, since this distribution is maxent among all the distributions with mean $\theta_{0}$ that are supported in $\mathbb{R}^{+}$. Consequently $\text{TP}_{\text{max}}=F_{0}(\ell_{X})=\int_{\ell_{X}}e^{-x/\theta_{0}}\text{d}x/\theta_{0}$.
@@
Second, the selection of the class $\mathcal{F}$ as a family of maximum entropy distributions simultaneously solves both the normalization and the selection bias problems [for a detailed discussion, see 22], maybe the two biggest concerns raised by physicists [17, 4, 23, 8] and philosophers [11, 12, 24] alike. On one side, considering more general classes of maxent distributions than uniform ones solves the normalization problem. At the same time, by considering the whole class $\mathcal{F}$, and not only the distribution induced by an estimate of $\theta_{0}$, from the observed value of $X$ in our universe, the selection bias problem is solved. Consider for instance the case when $\mathcal{X}=\mathbb{R}^{+}$ and $\mathcal{F}$ is the class of exponential distributions with expected value $\theta\in\mathbb{R}^{+}$. It can then be shown that the maximum likelihood estimator of $\theta_{0}$ is $\hat{\theta}_{0}=X$. But we still have no guarantee that $F_{0}(\ell_{X})=F(\ell_{X};\theta_{0})\leq F(\ell_{X};\hat{\theta}_{0})$. On the other hand, by (2), we know that $F_{0}(\ell_{X})\leq\text{TP}_{\text{max}}$. Thus, considering all possible values of $\theta$ removes the bias induced by the weak anthropic principle, since the class $\mathcal{F}$ includes $F(\cdot,\theta_{0})$ but is not reduced to it.
@@
Third, Algorithm 1 reveals that the output heavily depends on the input, namely the sample space $\mathcal{X}$ for $X$ and the constraints to impose on the family ${\cal F}$ of prior distributions $F$ of $X$. Moreover, Díaz-Pachón, Hössjer, and Marks proved a theorem (Appendix 3 of Díaz-Pachón et al. [22], summarized in Table 1 below, and explained in Remark 1) where they show that subtle changes in the input might produce extremely different outcomes. This can be seen, for instance, from Rows 4-5 of Table 1, when considering families $\cal F$ of distributions over $\mathcal{X}=\mathbb{R}$ that depend on a scale parameter only: if $0\in\ell_{X}$, it will provoke $\text{TP}_{\text{max}}=1$, regardless the size of the interval $\ell_{X}$ (therefore the level of tuning cannot be assessed), whereas $0\notin\ell_{X}$ with midpoint $x_{0}$ of $\ell_{X}$, will produce a small $\text{TP}_{\text{max}}$ provided the relative half size $\epsilon=|\ell_{X}|/(2x_{0})$ of the interval, is small. In the same direction, when $\mathcal{X}=\mathbb{R}^{+}$ ($\mathcal{=}\mathbb{R}$) and the form and scale (location and scale) family $\mathcal{F}$ is considered, $\text{TP}_{\text{max}}$ jumps from very small to 1, depending on whether the maximal signal-to-noise ratio $T$ of the prior distributions $F\in\mathcal{F}$ in Rows 2-3 (7-8) is bounded or not.
@@ 2 Mathematical Framework for Learning and Knowledge Acquisition
The No Free Lunch theorems [25, 26] assert that, on average, no search does better than a blind one, and therefore a guided search infuses information to the search problem. Active information was thus defined in order to measure the amount of information introduced by a programmer in order to reach a target, compared to a baseline distribution which is usually, but not necessarily, in maximum entropy [27, 28]. Formally, active information is defined as
@@
where the target $A\subset\Omega$ is a subset of a search space $\Omega$, whereas ${\mathbf{P}}$ and ${\mathbf{P}}_{0}$ are probability distributions on $\Omega$ that represent searches of the programmer, and blind search, respectively.
Since its inception, active information has been used in the measurement of bias for machine learning algorithms [29, 30, 31, 32], hypothesis testing [33, 34], statistical genetics [35, 36], bump hunting [37, 38], and estimation and correction of prevalence estimators of Covid-19 [39, 40], among others. In fact, active information can be used as a measure of FT if the search space $\Omega$ equals the sample space $\mathcal{X}$ of the physical parameter $X$, (4) is large for $A=\ell_{X}$, with ${\mathbf{P}}_{0}(A)=F_{0}(A)$ as in (1) and ${\mathbf{P}}(A)=\delta_{x_{0}}(A)$ a one-point distribution at $x_{0}$, the observed value of $X$ (or the midpoint of $\ell_{X}$) [10, 41].
Based also on active information, a mathematical formalization of the epistemological notions of learning and knowledge acquisition was recently developed by Hössjer et al. [42], where knowledge is usually defined as “justified true belief” [43, 44, 45]. This means that a subject or agent $S$ knows a proposition $p$ if
@@
Moreover, the inclusions are proper, since it is possible to have a belief in a false proposition so that such a belief does not constitute learning, and it is also possible to learn $p$ without getting to know $p$, if the true belief cannot be justified.
As for the mathematical formalization, belief is defined as a probability, as it is customary in Bayesian theory [see, e.g., 46, 47], whereas propositions are parameters that have a true value. More explicitly, suppose we want to learn whether a proposition $p$ is either true or false. A set of possible worlds $\Omega$ is defined, where one world $\omega_{0}$ represents the value of the parameter $\omega$, referred to as true world, whereas all other $\omega\in\Omega\setminus\{\omega_{0}\}$ are counterfactuals. The set $A$ of interest is made of the worlds in which proposition $p$ is true. An ignorant person assigns beliefs to every subset of $\Omega$ according to some initial distribution ${\mathbf{P}}_{0}$, whereas an agent $S$ with some data $D$ and discernment $\mathcal{G}$ (corresponding to a $\sigma$-algebra generated by the non-trivial events which data cannot discern into smaller events) updates his beliefs to ${\mathbf{P}}$, according to Bayes’s rule,
@@
I) Agent $S$ has learnt about proposition $p$ compared to an ignorant person, either when $p$ is true and the posterior belief ${\mathbf{P}}$ about $p$ is higher than the prior belief ${\mathbf{P}}_{0}$ about $p$, or when $p$ is false and the posterior belief about $p$ is smaller than the prior belief about $p$. II) The agent $S$ has fully learnt $p$ (regardless of the beliefs of the ignorant person) if the posterior belief ${\mathbf{P}}$ about $p_{1}$ is 1 (0) when $p_{1}$ is true (false).
@@
In particular, a subject in maximum state of ignorance is represented by a maximum entropy (maxent) distribution ${\mathbf{P}}_{0}$ over $\Omega$.
@@
    $S$ has learnt about $p$ compared to the ignorant person.
@@
    The true world $\omega_{0}$ is among the pool of possibilities for $S$ according to his posterior beliefs.
@@
Mathematically, the three requirements of knowledge acquisition for Part I of Definition 3, amount respectively to

### 01 Mar 2024 v2

@@ Abstract
Fine-tuning studies whether some physical parameters, or relevant ratios between them, are located within so-called life-permitting intervals of small probability outside of which carbon-based life would not be possible. Recent developments have found estimates of these probabilities that circumvent previous concerns of measurability and selection bias. However, the question remains if fine-tuning can indeed be known. Using a mathematization of the concepts of learning and knowledge acquisition, we argue that most examples that have been touted as fine-tuned cannot be formally assessed as such. Nevertheless, fine-tuning can be known when the physical parameter is seen as a random variable and it is supported in the nonnegative real line, provided the size of the life-permitting interval is small in relation to the observed value of the parameter.
Anthropic principle, Bayesian statistics, Analytical mathematics, Cosmological constant, Cosmological parameters
@@ 1 Introduction
Cosmological fine-tuning (FT) says that some physical parameters, some ratios between them, and some boundary conditions, must pertain to intervals of small probability to permit the existence of carbon-based life (Lewis & Barnes, 2016). FT started to spark the interest of the scientific community when Carter (1974) brought it to light calling it the anthropic principle, a term that, according to Lewis & Barnes (2016), was popularized and modified by Barrow & Tipler (1988). Since then, considerable effort has been made to determine the length of these so-called life-permitting intervals (LPIs), constituting the area where the scientific literature has primarily focused (see, e.g., Adams, 2019; Barnes, 2012; Davies, 1982). For this reason, some have linked FT with the small length of the intervals, with no regard for their probability (see, e.g., Helbig, 2023). This evokes the idea of “naturalness” in physics, where a good theory is assumed to present numbers that are close to unity, and departures from it point to ad-hoc explanations trying to fit observations to theory as Mercury epicycles in the Ptolemaic model (Hossenfelder, 2020). FT goes beyond naturalness, asserting that, according to our current theories, such unnaturalness is needed if carbon-based life is going to exist (Barnes, 2021). The existence of life sets a specification; i.e., a subset of possible outcomes of parameters for which life is permitted. These outcomes maximize a function $f$ that quantifies how specified outcomes are, and it is stochastically independent of the original random variable. In cosmological FT, this random variable is the physical parameter, whose value is assumed to be independent of the requisites for carbon-based life existence, and the LPI maximizes an indicator function $f$ that determines if life is possible or not. Therefore, the tuning of a physical constant for life is fine if and only if the probability of its LPI is small (Díaz-Pachón & Hössjer, 2022).
@@
Fine-tuning divides the scientific and philosophical communities between those who dismiss it as mere speculation (e.g., Hossenfelder, 2020; Colyvan et al., 2005; McGrew et al., 2001) and those who consider it a serious scientific endeavor (e.g., Davies, 2008; Lewis & Barnes, 2016; Tegmark, 2015). One of the main criticisms, even among those who see it as a legitimate question, is the lack of a valid probability distribution to impose over the parameters (Adams, 2019). In this direction, Hossenfelder asks: “[H]ow do you make a statement about probability without a probability distribution?” (Hossenfelder, 2020, p. 205). The problem is indeed daunting: it requires estimating the probability of the LPI using a biased sample of size 1 taken from an unknown distribution supported in an unknown space. More explicitly, and to establish some notation, let $\mathcal{X}$ be the possible values of a random physical parameter $X$. FT is about inferring the probability of its LPI $\ell_{X}$ when the only observation we possess is $x_{0}$, the observed value of $X$ in our universe; moreover, such an observation suffers from selection bias because that universe harbors life; and the only information about the probability distribution $F=F_{X}$ of $X$ is that $F$ is supported on $\mathcal{X}$ and $x_{0}\in\mathcal{X}$. Moreover, assuming Carter’s definition, it is just natural to assume that $F$ corresponds to a prior distribution of $X$ that maximizes the entropy (Carter, 1974).
@@
Past efforts to find the probabilities of LPIs use mainly uniform distributions over the set of possible values that constants might take (see, e.g., Sandora, 2019a; Barnes, 2019-2020; Tegmark et al., 2006). Nonetheless, this approach attests to the difficulty of the task, since the choice of the uniform distribution is based on Bernoulli’s Principle of Insufficient Reason (Bernoulli, 1713), which requires knowledge that the sample space is finite —a strong assumption that cannot be warranted for cosmological FT. This legitimate criticism of FT measurements came from philosophy, where it was called the normalization problem (Colyvan et al., 2005; McGrew et al., 2001; McGrew & McGrew, 2005; McGrew, 2018).
@@
Recently Díaz-Pachón, Hössjer, and Marks developed Algorithm 1 (below) to find an upper bound on the probability of the known LPI $\ell_{X}$ of a parameter $X$ (Díaz-Pachón et al., 2021, 2023). Algorithm 1 can be viewed as a statistical test, where the null hypothesis $H_{0}$ that tuning of $X$ is coarse is tested against the alternative hypothesis $H_{1}$ that $X$ is fine-tuned. The first two steps of Algorithm 1 ask for the input: a parameter space $\mathcal{X}$ (e.g., $\mathbb{N}$, $\mathbb{R}^{+}$, $\mathbb{R}$, $\mathbb{R}^{n}$, $\mathbb{C}$, or some subset thereof), and the constraints to impose on the prior distributions $F_{X}$ supported on $\mathcal{X}$ (e.g., if some event $B\subset\mathcal{X}$ has a known probability, if some of the moments of $X$ are finite, etc.). In the third step, a collection
@@
is obtained. In the fifth step, the outcome of Algorithm 1 is produced: if $\text{TP}_{\max}$ is small ($\leq\delta$ according to Definition 1), we reject the null hypothesis $H_{0}$ and conclude there is FT. If $\text{TP}_{\max}$ is not small, the algorithm is inconclusive. In this case, we cannot reject $H_{0}$. This does not mean that $X$ is not finely tuned though. Indeed, $H_{1}$ may still be true when $H_{0}$ is not rejected, since $F_{0}(\ell_{X})$ may be larger or smaller than $\delta$ when $\mathrm{TP}_{\max}>\delta$.
{algorithm}
Algorithm for testing fine-tuning
Input: Choose the set of possible values $\mathcal{X}$ of $X$.
Input: Define constraints on the distribution $F=F_{X}$ of $X$.
@@
Find the family =${\cal F}=\{F(\cdot;\theta);\theta\in\Theta\}$ of maxent distributions $F$ with support $\mathcal{X}$, subject to the constraints in Step 2, and assume ($H_{0}$) that the true distribution of $X$ satisfies $F_{0}=F(\cdot;\theta_{0})\in\mathcal{F}$ for some $\theta_{0}\in\Theta$.
@@
Find the maximum tuning probability $\mathrm{TP}_{\max}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\max\{F(\ell_{X});\,F\in{\cal F}\}$ over $\mathcal{F}$ for the life-permitting interval $\ell_{X}$.
@@
Output: If $\mathrm{TP}_{\max}$ is small ($\leq\delta$) there is fine-tuning (reject $H_{0}$); otherwise, the test is inconclusive (do not reject $H_{0}$).
@@
Algorithm 1 has some important properties. First, the upper bound $\mathrm{TP}_{\max}$ is sharp because if the family ${\cal F}$ is made of a single maxent distribution $F_{0}$, then $\text{TP}_{\max}=F_{0}(\ell_{X})$. For instance, assume that Step 1 selects $\mathcal{X}=\mathbb{R}^{+}$, and in Step 2 it is assumed that the first moment is finite and known (e.g., ${\mathbf{E}}_{0}X=\theta_{0}$). Then the family of Step 3 is made of a single distribution with density $F^{\prime}(x;\theta_{0})=f(x;\theta_{0})=e^{-x/\theta_{0}}/\theta_{0}$, the exponential with mean $\theta_{0}$, since this distribution is maxent among all the distributions with mean $\theta_{0}$ that are supported in $\mathbb{R}^{+}$. Consequently, $\text{TP}_{\text{max}}=F_{0}(\ell_{X})=\int_{\ell_{X}}e^{-x/\theta_{0}}\text{d}x/\theta_{0}$.
@@
Second, the selection of the class $\mathcal{F}$ as a family of maxent distributions simultaneously solves both the normalization and selection bias problems (for a detailed discussion, see Díaz-Pachón et al., 2023), maybe the two biggest concerns raised by physicists (Tegmark et al., 2006; Adams, 2019; Hossenfelder, 2021, 2020) and philosophers (Colyvan et al., 2005; McGrew et al., 2001; Bostrom, 2002) alike. On one side, considering more general classes of maxent distributions than uniform ones solves the normalization problem. At the same time, by considering the whole class $\mathcal{F}$, and not only the distribution induced by an estimate of $\theta_{0}$, from the observed value of $X$ in our universe, the selection bias problem is solved. Consider for instance the case when $\mathcal{X}=\mathbb{R}^{+}$ and $\mathcal{F}$ is the class of exponential distributions with expected value $\theta\in\mathbb{R}^{+}$. It can then be shown that the maximum likelihood estimator of $\theta_{0}$ is $\hat{\theta}_{0}=X$. But we still have no guarantee that $F_{0}(\ell_{X})=F(\ell_{X};\theta_{0})\leq F(\ell_{X};\hat{\theta}_{0})$. On the other hand, by (2) we know that $F_{0}(\ell_{X})\leq\text{TP}_{\text{max}}$. Thus, considering all possible values of $\theta$ removes the bias induced by the weak anthropic principle, since the class $\mathcal{F}$ includes $F(\cdot,\theta_{0})$ but is not reduced to it.
@@
Third, Algorithm 1 reveals that the output heavily depends on the input, namely the sample space $\mathcal{X}$ for $X$ and the constraints to impose on the family ${\cal F}$ of prior distributions $F$ of $X$. Moreover, Díaz-Pachón, Hössjer, and Marks proved a theorem (Appendix 3 of Díaz-Pachón et al. (2023), summarized in Table 1 below, and explained in Remark 1) where they show that subtle changes in the input might produce extremely different outcomes. This can be seen, for instance, from Rows 4-5 of Table 1, when considering families $\cal F$ of distributions over $\mathcal{X}=\mathbb{R}$ that depend on a scale parameter only: if $0\in\ell_{X}$, it will provoke $\text{TP}_{\text{max}}=1$, regardless the size of the interval $\ell_{X}$ (therefore the level of tuning cannot be assessed), whereas $0\notin\ell_{X}$ with midpoint $x_{0}$ of $\ell_{X}$, will produce a small $\text{TP}_{\text{max}}$, provided the relative half size $\epsilon=|\ell_{X}|/(2x_{0})$ of the interval is small. In the same direction, when $\mathcal{X}=\mathbb{R}^{+}$ ($\mathcal{=}\mathbb{R}$) and the form and scale (location and scale) family $\mathcal{F}$ is considered, $\text{TP}_{\text{max}}$ jumps from very small to 1, depending on whether the maximal signal-to-noise ratio $T$ of the prior distributions $F\in\mathcal{F}$ in Rows 2-3 (7-8) is bounded or not.
@@ 2 Mathematical Framework for Learning and Knowledge Acquisition
The No Free Lunch theorems (Wolpert & MacReady, 1995, 1997) assert that, on average, no search does better than a blind one, and therefore a guided search infuses information to the search problem. Active information was thus defined to measure the amount of information introduced by a programmer in order to reach a target, compared to a baseline distribution which is usually, but not necessarily, in maxent (Dembski & Marks II, 2009; Díaz-Pachón & Marks II, 2020). Formally, active information is defined as
@@
where the target $A\subset\Omega$ is a subset of a search space $\Omega$, whereas ${\mathbf{P}}$ and ${\mathbf{P}}_{0}$ are probability distributions on $\Omega$ that represent searches of the programmer and blind search, respectively.
Since its inception, active information has been used in the measurement of bias for machine learning algorithms (Montañez, 2017a, b; Montañez et al., 2019, 2021), hypothesis testing (Díaz-Pachón et al., 2020; Hom et al., 2023), statistical genetics (Thorvaldsen & Hössjer, 2023; Díaz-Pachón & Marks II, 2020), bump hunting (Díaz-Pachón et al., 2019; Liu et al., 2023), and estimation and correction of prevalence estimators of Covid-19 (Hössjer et al., 2023; Zhou et al., 2023), among others. In fact, active information can be used as a measure of FT if the search space $\Omega$ equals the sample space $\mathcal{X}$ of the physical parameter $X$, (4) is large for $A=\ell_{X}$, with ${\mathbf{P}}_{0}(A)=F_{0}(A)$ as in (1) and ${\mathbf{P}}(A)=\delta_{x_{0}}(A)$ a one-point distribution at $x_{0}$, the observed value of $X$ (or the midpoint of $\ell_{X}$) (Díaz-Pachón & Hössjer, 2022; Thorvaldsen & Hössjer, 2020).
Based also on active information, a mathematical formalization of the epistemological notions of learning and knowledge acquisition was recently developed by Hössjer et al. (2022), where knowledge is usually defined as “justified true belief” (Gettier, 1963; Ichikawa & Steup, 2018; Schwitzgebel, 2021). This means that a subject or agent $S$ knows a proposition $p$ if
@@
Moreover, the inclusions are proper, since it is possible to have a belief in a false proposition so that such a belief does not constitute learning; and it is also possible to learn $p$ without getting to know $p$, if the true belief cannot be justified.
As for the mathematical formalization, belief is defined as a probability, as it is customary in Bayesian theory (see, e.g., Berger, 2010; MacKay, 2003), whereas propositions are parameters that have a true value. More explicitly, suppose we want to learn whether a proposition $p$ is either true or false. A set of possible worlds $\Omega$ is defined, where one world $\omega_{0}$ represents the value of the parameter $\omega$, referred to as true world, whereas all other $\omega\in\Omega\setminus\{\omega_{0}\}$ are counterfactuals. The set $A$ of interest is made of the worlds in which proposition $p$ is true. An ignorant person assigns beliefs to every subset of $\Omega$ according to some initial distribution ${\mathbf{P}}_{0}$, whereas an agent $S$ with some data $D$ and discernment $\mathcal{G}$ (corresponding to a $\sigma$-algebra generated by the non-trivial events which data cannot discern into smaller events) updates his beliefs to ${\mathbf{P}}$, according to Bayes’s rule,
@@
I) Agent $S$ has learnt about proposition $p$, compared to an ignorant person, either when $p$ is true and the posterior belief ${\mathbf{P}}$ about $p$ is higher than the prior belief ${\mathbf{P}}_{0}$ about $p$, or when $p$ is false and the posterior belief about $p$ is smaller than the prior belief about $p$. II) The agent $S$ has fully learnt $p$ (regardless of the beliefs of the ignorant person) if the posterior belief ${\mathbf{P}}$ about $p_{1}$ is 1 (0) when $p_{1}$ is true (false).
@@
In particular, a subject in maximum state of ignorance is represented by a maxent distribution ${\mathbf{P}}_{0}$ over $\Omega$.
@@
    $S$ has learnt about $p$, compared to the ignorant person.
@@
    The true world $\omega_{0}$ is among the pool of possibilities for $S$, according to his posterior beliefs.
@@
Mathematically, the three requirements of knowledge acquisition for Part I of Definition 3 amount respectively to